OpenStudy (anonymous):

Which algebraic expression is equivalent to the expression below? (14x + 9) - 9x A 14x B 5x + 9 C 5x - 9 D 23x + 9

OpenStudy (midhun.madhu1987):

14x + 9 - 9x Combine the like terms

OpenStudy (anonymous):

@baking4lyf

OpenStudy (anonymous):

u know

OpenStudy (anonymous):

?

OpenStudy (midhun.madhu1987):

(14x - 9x) + 9 can u do it now?

OpenStudy (anonymous):

i think so

OpenStudy (anonymous):

Then if that i will go with D But if I were you I will wait for someone else to see if my answer is right

OpenStudy (anonymous):

Good Luck!

OpenStudy (midhun.madhu1987):

14-9 = ?

OpenStudy (anonymous):

sry 14-9=5

OpenStudy (anonymous):

@midhun.madhu1987

OpenStudy (midhun.madhu1987):

yes... so 14x - 9x = 5x. so which option is the answer?

OpenStudy (anonymous):

B right?

OpenStudy (anonymous):

@midhun.madhu1987

OpenStudy (midhun.madhu1987):

Yeah... Perfect :)

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

CORRECT ANSWERS YO